OpenStudy (anonymous):

Someone please show the steps how to do this I will medal!!! Factor each expression. Check your answer. w^2-7w-8

OpenStudy (ciarán95):

If you simply mean the expression w^2 - 7w -8, we can factorise as follows: w^2 - 7w -8 = (x - 8)(x + 1) (If you multiply this out you can see it gives us back the original expression).

OpenStudy (anonymous):

Thanx can you help with one more?

OpenStudy (ciarán95):

Sure i'll try and help you

OpenStudy (anonymous):

Factor each expression completly. 20w^2-45w+10

OpenStudy (ciarán95):

Here, we can make life easier by simplifying the expression first. So, if we divide everything in the expression by 5, we get 4w^2 - 9w +2. This make it easier to find the factors (of 4 and 2 now) to get the correct answer. So: 4w^2 - 9w +2 =(4w - 1)(w - 2) Lots of different methods to solving this, but the main thing to aim for is to get the middle value ('9' in the above case) from inserting factors of the 4 and 2 (4x1 and 2x1 here), although it can take a while to spot them. We can always check by multiplying out our answer and we should get the original expression.

OpenStudy (anonymous):

Thanks!!